# Universal Counterdiabatic Driving

### v1

@@ II.2 Noninteracting models
Figure 3: Approximation of $1/x$ by Chebyshev polynomials, where $x=\omega/\Omega$. The inset shows the same plot on a linear scale. In each case, the optimal $\mu(\ell)$ that is obtained by maximizing fidelity for the TFI model annealing problem is used, with fixed $\Omega(\ell)=\Omega_{max}$.
@@ II.3 Interacting Models & Asymptotes of the Spectral Function
For single-particle models, where there are no excitations above a certain maximum, this method is guaranteed to converge since for constant $\Omega(\ell)$, clearly $\lim_{\ell\to\infty}\mu(\ell)=0$. Thus, the interval $[\mu(\ell),\Omega]\to_{l\to\infty}[0,\Omega]$ such that the universal protocol will asymptotically suppress more and more low-energy excitations with increasing $\ell$ without exciting any high-frequency modes. In general, the convergence of the universal protocol with $\ell$ is guaranteed as long as $\Omega(\ell)$ grows slower than $\ell/\log(\ell)$. In this case we still have $\lim_{\ell\to\infty}\mu(\ell)=0$.
@@
We see that in order to control the UV error of the cost function we should choose $\Omega(\ell)$ to be at least as the bound given by Eq. (19). As a result using Eq. (17) and ignoring for simplicity logarithmic corrections we arrive at the scaling
@@
Summarizing our discussion here we arrive at the most surprising conclusion in this work, tying the complexity of local CD protocols to the high frequency properties of the system. This happens despite the fact that when constructing counterdiabatic protocols, we are typically interested in suppressing the low energy excitations in the system (which set the norm of the exact AGP). The same low energy excitations determine the minimal protocol time in conventional quantum annealing schemes. For example, if we analyze driving across a critical point, such low-energy excitations get “frozen out” in the Kibble-Zurek mechanism Zurek (1985); Polkovnikov (2005); Dziarmaga (2005); Zurek et al. (2005). The scaling (20) states that contrary to these expectations, performance of the local CD protocol – at least in the Krylov space – is tied to the high frequency response of the system. This suggests there is some fundamental connection between the short- and long-time dynamics as alluded to earlier.
@@ III.1 Transverse Field Ising Model with Modulated Fields
with $h_{i}$ drawn uniformly between 00 and $h$ with $h=J=1$. This will increase the size of the blocks in momentum space after the Jordan-Wigner transformation, but the blocks remain small enough that the dynamics may be computed via exact diagonalization.
@@ III.2 Ising Model with Next-Nearest-Neighbor Interactions
With the scaling of $\zeta(\ell)$ from Eq. (17), this linear growth of $\Omega(\ell)$ gives $\mu(\ell)\sim\log\ell$. So the infrared cutoff, instead of decreasing, slowly increases with $\ell$. We therefore conclude that in the thermodynamic limit $N\to\infty$ the universal protocol, and more generally the Krylov space variational ansatz for the AGP does not converge. Notice though that the universal protocol still leads to a significant improvement of fidelity at finite $\ell$.
@@ III.3 XXZ Model
Figure 6: For the XXZ model with $N=18$, annealing from the Heisenberg to anisotropy dominated point, the optimal $\Omega(\ell)$ is obtained numerically by maximizing the fidelity as before. The dashed line is the best fit to a power-law of the data, showing a slope 0.52 which is very close to the Section II.3 prediction of 0.5 with a logarithmic correction. The asymptotic behavior of the fidelity, inset, is hard to predict for the finite sized systems accessible. But it does not appear to plateau.
IV Summary & Outlook
@@
In this work we minimize this distance using a fixed Chebyshev polynomial approximation of $1/\omega$ in a finite window $[\zeta(\ell),1]$, where $\zeta(\ell)\sim\log(\ell)/\ell$ and $\ell$ is the order of the polynomial. Then we use an overall energy scale $\Omega(\ell)$, which rescales this interval to $[\mu(\ell),\Omega(\ell)]$ where $\mu(\ell)=\zeta(\ell)\Omega(\ell)$, as the single remaining variational parameter. We find that its asymptotic dependence on $\ell$ is determined by the scaling of the high-frequency tail of the spectral function. In particular, if these tails decay faster than exponentially with frequency, the universal protocol is guaranteed to improve with $\ell$ even in the thermodynamic limit. However, in generic situations, where these tails are expected to be exponential, we find that the scale $\Omega(\ell)$ needs to grow linearly with $\ell$ so that the lower cutoff $\mu(\ell)$ saturates at a finite value. This saturation prevents an approximate CD protocol from continuously improving with $\ell$ in the thermodynamic limit.
These protocols perform comparably with variational approaches that require knowledge of the Hamiltonian beforehand. Along the way, we discover a very interesting connection between the low- and high-frequency response of a system.
There are several further improvements to this method that might be made. For example, the high-frequency tail of the spectral function usually converges quickly with the system size, such that one could compute this efficiently for small systems and extrapolate to the thermodynamic limit. Then one can use this extrapolated spectral function as a weight function in the polynomial fitting of $1/\omega$ and additionally improve convergence of the approximate CD protocol Finžgar et al. (2025).
Interestingly, the results of this work show that one can search for optimal annealing paths amenable to local CD driving by minimizing high-frequency tails of the spectral function and hence minimizing high-frequency noise and dissipation. These paths should be most suitable for fast adiabatic state preparation. This finding suggests existence of some fundamental connections between short- and long-time response of interacting systems and needs further exploration. Lastly, this method might be fruitfully applied to practical quantum annealing where the problem Hamiltonian, or a model for the disorder/noise, is not exactly known.
@@ Acknowledgments
The authors thank Dries Sels for many beneficial discussions, and John Martyn for information on rigorous results in polynomial approximation. This work was supported by the AFOSR Grant FA9550-21-1-0342. The exact code used is available online  [^3] See https://github.com/smorawetz/CD-extra-controls.git.
@@ Appendix A Extra factor of $\omega$ in cost function
[…]
@@
[…]
@@ Appendix C Numerical evidence of spectral function Gaussian decay in XXZ Model
Using a recently developed Julia package leveraging the Pauli string structure of spin-1/2 models Loizeau et al. (2024), we can compute these Lanczos coefficients for very large systems. Although $\alpha=2$ has been observed several times at the Heisenberg point, the growth at other nearby points has not been studied. We reparametrize the XXZ model as

### v2

@@ II.2 Noninteracting models
Figure 3: Approximation of $1/x$ by Chebyshev polynomials, where $x=\omega/\Omega$. The inset shows the same plot on a log scale. In each case, the optimal $\mu(\ell)$ that is obtained by maximizing fidelity for the TFI model annealing problem is used, with fixed $\Omega(\ell)=\Omega_{max}$.
@@ II.3 Interacting Models & Asymptotes of the Spectral Function
For single-particle models, where there are no excitations above a certain maximum, this method is guaranteed to converge since for constant $\Omega(\ell)$, clearly $\lim_{\ell\to\infty}\mu(\ell)=0$. Thus, the interval $\mu(\ell),\Omega]\xrightarrow[]{\ell\to\infty}[0,\Omega]$ such that the universal protocol will asymptotically suppress more and more low-energy excitations with increasing $\ell$ without exciting any high-frequency modes. In general, the convergence of the universal protocol with $\ell$ is guaranteed as long as $\Omega(\ell)$ grows slower than $\ell/\log(\ell)$. In this case we still have $\lim_{\ell\to\infty}\mu(\ell)=0$.
@@
We see that in order to control the UV error of the cost function we should choose $\Omega(\ell)$ to be at least as large as the bound given by Eq. (19). As a result, using Eq. (17) and ignoring for simplicity logarithmic corrections we arrive at the scaling
@@
Summarizing our discussion here we arrive at the most surprising conclusion in this work, tying the complexity of local CD protocols to the high frequency properties of the system. This happens despite the fact that when constructing counterdiabatic protocols, we are typically interested in suppressing the low energy excitations in the system (which set the norm of the exact AGP in Eq. (3)). The same low energy excitations determine the minimal protocol time in conventional quantum annealing schemes. For example, if we analyze driving across a critical point, such low-energy excitations get “frozen out” in the Kibble-Zurek mechanism Zurek (1985); Polkovnikov (2005); Dziarmaga (2005); Zurek et al. (2005). The scaling (20) states that contrary to these expectations, performance of the local CD protocol – at least in the Krylov space – is tied to the high frequency response of the system. This suggests there is some fundamental connection between the short- and long-time dynamics as alluded to earlier. This point will be further developed in Section IV.
@@ III.1 Transverse Field Ising Model with Modulated Fields
with $h_{i}$ drawn uniformly between 0 and $h$ with $h=J=1$. This will increase the size of the blocks in momentum space after the Jordan-Wigner transformation, but the blocks remain small enough that the dynamics may be computed via exact diagonalization.
@@ III.2 Ising Model with Next-Nearest-Neighbor Interactions
With the scaling of $\zeta(\ell)$ from Eq. (17), this linear growth of $\Omega(\ell)$ gives $\mu(\ell)\sim\log\ell$. Therefore the infrared cutoff, instead of decreasing, slowly increases with $\ell$. We therefore conclude that in the thermodynamic limit $N\to\infty$ the universal protocol, and more generally the Krylov space variational ansatz for the AGP does not converge. Additionally, the universal protocol still leads to a significant improvement of fidelity at finite $\ell$.
@@ III.3 XXZ Model
Figure 6: For the XXZ model with $N=18$, annealing from the Heisenberg to Ising point, the optimal $\Omega(\ell)$ is obtained numerically by maximizing the fidelity as before. The dashed line is the best fit to a power-law of the data, showing a slope 0.52 which is very close to the Section II.3 prediction of 0.5 with a logarithmic correction. The asymptotic behavior of the fidelity, inset, is hard to predict for the finite sized systems accessible. But it does not appear to plateau.
IV Extracting hydrodynamic tails from Lanczos coefficients
Across a wide variety of interacting systems, the long-time dynamics may often be described in some universal way, such as e.g. charge or energy diffusion in systems with corresponding conservation laws (see Ref. Hohenberg and Halperin, 1977 for a review). Recently, there has been extensive research on so-called generalized hydrodynamics where analogous universal long-time dynamics are seen in integrable models. Doyon (2020); Bertini et al. (2016); Malvania et al. (2021)
Another area of active investigation is operator spreading. In particular, in a series of works it was argued that short-time dynamics of a quantum system in Krylov space is universal and can be characterized by the Krylov complexity Parker et al. (2019); Avdoshkin and Dymarsky (2020); Rabinovici et al. (2021). In this approach time evolution of an observable in quantum (or classical) Hamiltonian system can be thought of as a one-dimensional hopping problem in the Krylov operator space formed from nested commutators of the operator describing this observable and the Hamiltonian like those entering Eq. (6). The hopping amplitudes are then given by the so called Lanczos coefficients appear in the Gram-Schmidt orthonormalization of these commutators (see e.g. Ref. Loizeau et al., 2025a for details). Formally all Lanczos coefficients can be found by analyzing a short time expansion of local operators in the Heisenberg representation. Finding if there are any connections between this short time evolution and long time hydrodynamic tails is an open and long-standing problem.
Unexpectedly, our work provides an insight into this. In particular, for systems whose UV spectral function tails decay faster than exponentially (corresponding to sublinear growth of Lanczos coefficients) using the universal AGP we can reconstruct the hydrodynamic tails of the spectral function, and hence the long-time asymptotic behavior of the correlation functions. Conversely, in systems with linear Lanczos coefficient growth, corresponding to exponential tails of the spectral function, the hydrodynamic tails could be extracted from the universal AGP only when the number of the Lanczos coefficients exceeds the system size. While we do not claim that there is no other method which can reveal such long time information from the Lanczos coefficients, it seems that the UV tails of the spectral function play a similar fundamental role in accessing this low frequency information as they do in constructing the universal CD protocol.
Now let us briefly outline how one can extract the hydrodynamic tails from the universal CD protocol. If we regard $\partial_{\lambda}H$ as a perturbation arising from changing $\lambda$, then the commutators $\hat{\mathcal{L}}^{2k-1}\partial_{\lambda}H$ which appear in the approximate AGP simply correspond to a short-time expansion of the evolution of this perturbation in the interaction picture. We can combine these with the coefficients $\alpha_{k}$ or $\beta_{k}$, which are agnostic to any physics, to compute a universal AGP. In this way we can determine the AGP norm (or fidelity susceptibility) using only these universal coefficients and the nested commutator norms (which are uniquely related to the Lanczos coefficients). In turn from Eq. (3) we see that the scaling of the fidelity susceptibility with $\mu$ is directly related to the low-frequency scaling of the spectral function (see also Refs. Kim and Polkovnikov, 2024; Pandey et al., 2020b). This scheme thus works as long as by increasing the number of terms in the universal protocol we can effectively access lower values of $\mu$. The central result of our work is that $\mu(\ell)$ decreases in the thermodynamic limit for systems with high frequency tails decaying faster than exponentially.
Let us now illustrate these considerations using the the example of the XXZ chain discussed in Section III.3 of the main text. There, since $\mu\sim 1/\sqrt{\ell}$ we expect that as $\ell$ is increased, the universal AGP should approach the true AGP. If the dynamics are diffusive as expected for the XXZ chain with $\Delta>1$ Gopalakrishnan and Vasseur (2019), then $\Phi_{\lambda}(\omega)\sim 1/\sqrt{\omega}$ Sels and Polkovnikov (2023b) and for large enough $\ell$ according Eq. (3)) $\chi_{\lambda}$ should scale with $\mu^{-3/2}$. For the universal AGP, we can compute this directly from the definition in Eq. (7):
$\|A_{\lambda}^{(\ell)}\|^{2}$ $=\sum_{m,n}\left(\sum_{k}\alpha_{k}(\mu,\Omega)\ \omega_{mn}^{2k-1}\right)^{2}|\langle m|\partial_{\lambda}H|n\rangle|^{2}$ (24)
$=\int_{-\infty}^{\infty}d\omega\left(\alpha_{k}(\mu,\Omega)\ \omega^{2k-1}\right)^{2}\Phi_{\lambda}(\omega)$ (25)
Here, we have written $\mu$ and $\Omega$ as arguments of $\alpha_{k}$ explicitly to highlight that they alone determine the coefficients. Then, we can compare numerically the norms of the approximate and exact AGPs for each $\mu$. This is shown in Figure 7. From this, it appears that the universal AGP is able to capture the diffusive hydrodynamics of the model, despite the only input about the model being information present in a short time expansion. We expect that the scaling will work even better if instead of exact diagonalization one uses other numerical tools allowing access to short time dynamics (and hence the Lanczos coefficients) in larger systems.
Finally let us point that in the generic models (say a non-integrable Ising model discussed here) increasing the number of terms $\ell$ in the universal protocol does not lead to decrease of $\mu(\ell)$ as long as $\ell\ll L$ [^3] We mention again that in the parallel paper Finžgar et al. (2025) it was argued that if one uses the information about high frequency tail of the spectral function one can make $\mu~\sim 1/\log(\ell)$ such that it slowly decreases with $\ell$. This implies that we cannot access the low $\mu$ asymptote of $\chi_{\lambda}$ and hence of the spectral function unless $\ell$ exceeds system size $L$ and the upper cutoff $\Omega$ stops growing with $\ell$.
Figure 7: On the left hand side, the exact AGP norm, computed via Eq. (3) is compared with that from the universal AGP in Eq. (24). The dashed lines show the expected asymptotic scaling for diffusion. The numerics are for the $N=18$ XXZ chain at $\lambda=3/8$, corresponding to $\Delta=1.2$ in standard parametrization of XXZ model. Minor finite size effects may play a role below $\log_{10}\mu\approx-0.3$ since the largest operators in the commutator expansion reach the boundary modifying the commutator norms compared to thermodynamic limit. On the right hand side, the exact spectral function for this model, averaged over several nearby values of $\lambda$, is plotted for different system sizes. The regime of $\omega$ at which diffusion is present slowly increases with system size and is expected to extend all the way to zero frequency in the thermodynamic limit. For $\mu$ within this region, we expect the scaling $\|A_{\lambda}\|^{2}\sim\mu^{-3/2}$.
V Summary & Outlook
@@
In this work we do this with a fixed Chebyshev polynomial approximation of $1/\omega$ in a finite window $\zeta(\ell),1]$, where $\zeta(\ell)\sim\log(\ell)/\ell$ and $\ell$ is the order of the polynomial. Then we use an overall energy scale $\Omega(\ell)$, which rescales this interval to $\mu(\ell),\Omega(\ell)]$ where $\mu(\ell)=\zeta(\ell)\Omega(\ell)$, as the single remaining variational parameter. We find that its asymptotic dependence on $\ell$ is determined by the scaling of the high-frequency tail of the spectral function. In particular, if these tails decay faster than exponentially with frequency, the universal protocol is guaranteed to improve with $\ell$ even in the thermodynamic limit. However, in generic situations, where these tails are expected to be exponential, we find that the scale $\Omega(\ell)$ needs to grow linearly with $\ell$ so that the lower cutoff $\mu(\ell)$ saturates at a finite value. This saturation prevents an approximate CD protocol from continuously improving with $\ell$ in the thermodynamic limit.
These protocols perform comparably with variational approaches that require knowledge of the Hamiltonian beforehand. There are several further improvements to this method that might be made. For example, the high-frequency tail of the spectral function usually converges quickly with the system size, such that one could compute this efficiently for small systems and extrapolate to the thermodynamic limit. Then one can use this extrapolated spectral function as a weight function in the polynomial fitting of $1/\omega$ and additionally improve convergence of the approximate CD protocol Finžgar et al. (2025).
Interestingly, the results of this work show that one can search for optimal annealing paths amenable to local CD driving by minimizing high-frequency tails of the spectral function and hence minimizing high-frequency noise and dissipation. These paths should be most suitable for fast adiabatic state preparation. This finding suggests the existence of some fundamental connections between short- and long-time response of interacting systems and needs further exploration. Lastly, this method might be fruitfully applied to practical quantum annealing where the problem Hamiltonian, or a model for the disorder/noise, is not exactly known.
@@ Acknowledgments
The authors thank Dries Sels for many beneficial discussions, as well as Hyeongjin Kim and John Martyn. This work was supported by the AFOSR Grant FA9550-21-1-0342 and the NSF Grant DMR-2412542. The exact code used is available online  [^4] See https://github.com/smorawetz/CD-extra-controls.git.
@@ Appendix A Extra factor of $\omega$ in cost function
$S_{\lambda}=\min_{\beta_{k}}\int d\omega\left(\sum_{k}\beta_{k}T_{2k}(\omega)+g(\omega/\mu)\right)^{2}\Phi_{\lambda}(\omega)$ (26)
@@
$\tilde{S}_{\ell}(\zeta)=\int_{\zeta}^{1}dx\left(\sum_{k=1}^{\ell}\beta_{k}T_{2k}(x)+1\right)^{2},$ (27)
@@ Appendix C Numerical evidence of spectral function Gaussian decay in XXZ Model
Using a recently developed Julia package leveraging the Pauli string structure of spin-1/2 models Loizeau et al. (2025b), we can compute these Lanczos coefficients for very large systems. Although $\alpha=2$ has been observed several times at the Heisenberg point, the growth at other nearby points has not been studied. We reparametrize the XXZ model as